# Governing Dynamic Capabilities: Cryptographic Binding and Reproducibility Verification for AI Agent Tool Use

### 15 Mar 2026 v1

@@ 8.1 Scope of Contribution
Our contribution is a governance framework, not novel cryptographic primitives. The individual building blocks—X.509 certificates, SHA-256, Ed25519, hash chains—are standard. The novelty lies in (a) identifying the capability-context separation as the principled boundary for agent governance, (b) showing that composing standard mechanisms along this boundary enforces G1–G3 with negligible overhead, and (c) demonstrating this is practically achievable. We view this as analogous to Certificate Transparency [16], which composes Merkle trees and gossip protocols—neither novel individually—into infrastructure that fundamentally changed how the web PKI operates. The security proofs are intentionally straightforward reductions to standard assumptions (collision resistance, EUF-CMA); the contribution is in what to build, not in new ways to prove things.
@@ 8.2 Capability Binding Limitations
Capability-context boundary. Our sharp boundary has edge cases: a “code execution” tool can effectively acquire new capabilities through composition (e.g., downloading and running arbitrary scripts). Our framework guarantees that the registered capability set has not changed, but the effective attack surface within a fixed capability envelope can still be large. Complementary approaches (sandboxing, taint tracking, output filtering) are needed to govern behavior within the envelope. We view capability binding and intra-capability governance as orthogonal layers: the former ensures the attack surface is known; the latter constrains behavior within it.
@@
Re-signing frequency. Mitigated through wildcard tool categories, short-lived auto-renewable certificates, and the empirical observation that tool configurations change far less frequently than user context.
8.3 Reproducibility Limitations and Graceful Degradation
Reproducibility is a design constraint, not a universal guarantee. Our framework makes explicit what is often implicit: non-reproducible agent behavior is non-verifiable. This is not a limitation unique to our approach—it applies to any verification scheme not based on trusted hardware. For high-assurance deployments, this means agents should be configured for deterministic output (structured formats, temperature==0, fixed seeds). Our empirical results (§7) show that code-generating agents already satisfy this constraint perfectly, and the separation between same-model and cross-model distributions ($11.5\times$) provides robust detection even for prose tasks.
Graceful degradation under partial G2. Our multi-provider experiment (§7, Experiment 5c) reveals that model determinism varies by over an order of magnitude: from CharMatch==0.987 (Haiku 4.5) to 0.169 (Kimi K2.5). This raises a practical question: what security guarantees hold when some agents use low-determinism models?
When G2 is unavailable (reproducibility commitment $\mathit{none}$), the framework degrades to G1+G3 only. Table 12 characterizes the precise security impact.
@@ 8.3 Reproducibility Limitations and Graceful Degradation
G1+G3 alone protect against 9 of 12 attack types. The three attacks that require G2—model substitution (ATK-2), cascading injection via model compromise (ATK-10/11 when the injection occurs through model replacement rather than tool replacement), and output exfiltration via model tampering (ATK-12)—share a common structure: the attacker operates within a valid capability envelope by substituting the computational process itself. Signing proves attribution; only G2 proves computational fidelity.
Chain-level degradation. The Chain Verifiability Theorem (§4.8) establishes a stronger result: a G2==none agent at an interior position breaks behavioral verification for all downstream nodes. This means the true cost of deploying a low-determinism model at a coordinator position is not merely that the coordinator itself is unverifiable, but that the entire sub-tree it governs loses G2 protection. The framework’s three commitment levels (full/statistical/none) are thus not equivalent—their security implications depend critically on chain position.
Statistical vs. computational security. Token-level reproducibility provides statistical rather than computational security. However, the Bounded Divergence Theorem (§4.7) shows this gap is operationally small: sufficient replay verification transforms statistical detection into a provable safety certificate. An adversary who fine-tunes a model to pass $n$ diverse replay checks has produced a model that is $(n,\epsilon)$-indistinguishable from the original, with $\epsilon$ decreasing exponentially in $n$. Future work will explore semantic similarity (embedding cosine distance) as a verification metric, which may provide tighter separation bounds and is robust to paraphrase-level variation.
Provider cooperation. Replay verification requires access to the same model version. Organizations should negotiate replay rights as part of service agreements. Deterministic inference providers [6] eliminate this dependency entirely.
@@ 8.4 Trust Infrastructure Limitations
Non-agent node compromise. We assume non-agent nodes (CAs, humans, organizations) are uncompromised. Standard PKI mitigations apply: Certificate Transparency logs, hardware trust anchors (TPM 2.0), and multi-party issuance protocols. Ledger operator trust. The current single-operator ledger relies on bilateral signing for tamper evidence. A multi-operator design with consensus is possible but introduces latency overhead. Application-layer data flow. Our framework governs agent identity and capability, not the content of data flows. Complementary approaches (data-flow taint tracking, differential privacy) address this orthogonal concern.
@@ 8.5 Future Work
Hardware-attested agent identity (TPM 2.0, TEEs) for stronger G2 guarantees without provider cooperation. Embedding-based semantic similarity for reproducibility verification with formal threshold analysis. Cross-protocol capability governance beyond MCP/A2A. AI-assisted certificate issuance to reduce operational burden. Dynamic trust calibration from ledger history, enabling trust scores to evolve based on observed agent behavior.
@@ 9 Related Work
Verifiable Execution. BAID [2] provides zkVM-based verification ($\sim$}120 s per proof). DiFR [1] demonstrates inference near-determinism. EigenAI [6] achieves 100% deterministic inference. Attestable Audits [7] uses TEEs.
@@
Agent Protocols. A2A [8] provides agent communication. MCP enables agent-tool interaction with OAuth 2.1.
@@ 10 Conclusion
AI agents increasingly acquire capabilities at runtime, yet no existing framework enforces the three governance requirements—capability integrity, behavioral verifiability, and interaction auditability—that a trustworthy agent ecosystem demands.
We addressed this through the capability-context separation principle and three mechanisms: capability-bound certificates (G1), reproducibility commitments (G2), and a verifiable interaction ledger (G3). We formalized nine security properties (including chain verifiability: behavioral verification is a chain property broken by any unverifiable interior agent), implemented the framework in Rust with 97 µs verification and <<1 ns capability binding overhead (${\sim}1{,}200{,}000\times$ faster than BAID), validated reproducibility across 9 models from 7 providers ($F_{1}=0.876$ multi-provider, $F_{1}=0.990$ single-provider), and verified 12 attack scenarios—including runtime behavioral attacks validated against NVIDIA Nemotron-AIQ traces—with specific, traceable detection mechanisms. An end-to-end evaluation deploying the full governance stack across a 5-to-20-agent pipeline with real LLM calls confirms 0.12% overhead, sub-linear per-agent scaling, and detection of all in-situ attacks with zero false positives—while three baselines (no governance, OAuth 2.1, OpenTelemetry) detect none.
Our framework fills the stateful, capability-aware governance gap between identity platforms and runtime gateways. As the EU AI Act’s high-risk provisions take effect and NIST develops agent-specific standards, the need for cryptographic infrastructure enforcing G1–G3 will only grow.
We release the core mechanisms as an open-source SDK (agent-trust-kit) to enable ecosystem adoption.

### 19 Mar 2026 v2

@@ 8.1 Scope of Contribution
Our contribution is a governance architecture, not new cryptographic primitives. The building blocks (digital signatures, hash chains, replay verification) are individually well-known; the contribution lies in identifying the capability-context separation as a principled governance boundary, proving structural constraints (Chain Verifiability, Bounded Divergence) that arise from this architecture, and defining abstract governance contracts (G1–G3) that hold across fundamentally different cryptographic backends.
Why the composition is non-obvious. A counterfactual test: given Ed25519, SHA-256, and hash chains, would a security engineer independently arrive at G1–G3? The evidence suggests not. The MCP ecosystem has grown to over 1,000 tool servers with OAuth 2.1 as its security mechanism, yet no deployment detects capability drift (G1 gap). OpenTelemetry is the industry standard for agent observability, yet the Nemotron-AIQ dataset [46] demonstrates that 10,796 OTEL traces provide no tamper evidence (G3 gap). BAID [2] constructs a zkVM over the full transformer pass, yet cannot detect whether the agent’s tool set has changed (G1 gap). Each existing approach addresses a fragment; none recognizes the capability-context separation as the organizing principle, and none arrives at the three-requirement conjunction that Table 1 shows is necessary and sufficient. The composition and the boundary identification are the contributions, validated by two instantiations (basic and enhanced) satisfying the same nine security properties under standard assumptions (CR, EUF-CMA; $q$-SDH, DDH, knowledge soundness).
8.2 The Agent Governance Trilemma
A fundamental question underlies the design of any agent governance framework: why can’t we simply have agents that are maximally capable, maximally fast, and maximally secure? We observe that this is not merely a current engineering limitation, but a structural constraint rooted in computability theory.
Observation 1 (Agent Governance Trilemma, informal).
For agent systems operating under a fixed computational budget, three desiderata—capability, performance, and security—cannot be simultaneously maximized. Formally:
  1. 1.
    High Capability + High Security $\Rightarrow$ Performance $\leq 1/2$. If an agent’s tool set is Turing-complete (unbounded capability) and full behavioral verification is required (high security), then verification costs at least as much as the original execution, so at most half the computational budget produces useful work.
  2. 2.
    Bounded Capability + High Security $\Rightarrow$ Performance $\to 1$. If the tool set is non-Turing-complete (bounded capability), verification reduces to lookup in a pre-computed table of valid input-output pairs, with $O(1)$ per-query overhead.
  3. 3.
    High Capability + High Performance $\Rightarrow$ Security degrades. If the agent is Turing-complete and most of the budget goes to useful work, insufficient resources remain for verification, and the verifier cannot detect all semantic deviations.
Proof sketch for (1). Behavioral verification (G2) requires determining whether the agent’s output $y$ matches what its declared model $M$ would produce on input $x$. For general computations, the only universal method is re-execution: compute $M(x)$ and compare. No sub-linear shortcut exists for arbitrary $M$, because $M$’s output depends on the full complexity of its parameters and computation—the function $M$ may be incompressible (its Kolmogorov complexity approaches its execution cost). Therefore $W_{\mathrm{verify}}\geq W_{\mathrm{task}}$, giving performance ratio $P=W_{\mathrm{task}}/(W_{\mathrm{task}}+W_{\mathrm{verify}})\leq 1/2$. For Byzantine-fault-tolerant verification with $k$ independent verifiers, $P\leq 1/(k+1)$.
Proof sketch for (2). When each tool $t_{i}$ has a finite input space $\mathcal{I}_{i}$ and output space $\mathcal{O}_{i}$, the set of valid $(x,t_{i}(x))$ pairs is bounded by $\sum_{i}|\mathcal{I}_{i}|$. A hash table of valid pairs can be pre-computed offline and queried in $O(1)$ per verification. Since the pre-computation is amortized over all future queries, the marginal verification cost is constant, and $P\to 1$ as the number of task executions grows.
Proof sketch for (3). This follows from Rice’s theorem: “the output of agent $A$ is safe” is a non-trivial semantic property of $A$’s program, and no algorithm can decide all such properties for Turing-complete programs. When the verification budget $W_{\mathrm{verify}}<W_{\mathrm{task}}$, the verifier cannot re-execute all computations. Sampling-based verification (checking $n$ out of $N$ outputs) provides statistical guarantees bounded by our Bounded Divergence Theorem ($\epsilon\leq\ln(1/\alpha)/n$), but the adversary can always deviate on the NnN-n unchecked outputs. Static analysis of agent programs is undecidable in the general case.
How TEE sidesteps the trilemma. Hardware TEE replaces computational verification with hardware attestation, shifting the tradeoff from the compute domain to the trust assumption domain (reliance on hardware vendor integrity and supply chain security). TEE does not break the trilemma but trades compute cost for trust assumptions, and satisfies only G2—a TEE-attested agent still requires G1 and G3.
Analogy to established impossibility results. The trilemma is structurally analogous to the CAP theorem [48] and the blockchain trilemma; we conjecture that a formal proof reducing to standard computability results is possible (§8.8).
8.3 Governance Levels: Operationalizing the Trilemma
The trilemma motivates three governance levels (Level 3: compile-time safety with bounded tools; Level 2: sampled verification for enterprise deployments; Level 1: post-hoc traceability for low-risk use cases), each corresponding to a different position in the capability-performance-security space. Full specifications including recommended crypto backends are given in Appendix B. The key insight is that every deployment’s security posture is legible: the governance level is declared in the agent’s certificate and recorded in the ledger, making the tradeoff explicit to auditors and downstream agents.
8.4 Capability Boundedness and Meta-Tools
The trilemma’s capability dimension distinguishes two tool classes. Bounded capabilities (web_search, read_file) have finite or structurally constrained input-output spaces; G1 fully characterizes the envelope and verification is efficient. Unbounded capabilities (code_execution, shell_access) are Turing-complete: G1 correctly records that the agent has the capability, but the envelope is formally declared yet semantically unbounded—knowing an agent can execute code does not predict what code it will execute.
For unbounded tools, governance either restricts to bounded mode (sandboxed execution, Level 3) or accepts unbounded mode with runtime monitoring ($P\leq 1/2$ per the trilemma, as the monitor consumes at least as much compute as the governed agent). G1–G3 provides the infrastructure for both: G1 declares boundedness, G2 verifies behavioral fidelity within the envelope, and G3 records all interactions for post-hoc audit.
@@ 8.5 Capability Binding Limitations
Re-signing frequency. Mitigated through wildcard tool categories (web_search@ˆ2.0) and short-lived auto-renewable certificates. A typical deployment (10 tools, weekly updates) incurs ${\sim}520$ certificate operations per year at <<100 µs each—operationally negligible.
8.6 Graceful Degradation under Partial G2
When G2 is unavailable (reproducibility commitment $\mathit{none}$), the framework degrades to G1+G3 only. Table 16 summarizes the security impact.
@@ 8.6 Graceful Degradation under Partial G2
G1+G3 alone protect against 9 of 12 attack types. The attacks requiring G2—model substitution (ATK-2), cascading injection via model compromise (ATK-10/11), and output exfiltration (ATK-12)—share a common structure: the attacker operates within a valid capability envelope by substituting the computational process itself.
Chain-level degradation. A G2==none agent at an interior position breaks behavioral verification for all downstream nodes (Chain Verifiability Theorem, §4.9); the deployment principles therein provide prescriptive architectural guidance.
Statistical vs. computational security. Token-level reproducibility provides statistical rather than computational security, but the Bounded Divergence Theorem (§4.8) bounds this gap: an adversary passing $n$ replay checks has produced a model $(n,\epsilon)$-indistinguishable from the original, with $\epsilon\leq\ln(1/\alpha)/n$.
Verification metric limitations. Current textual similarity metrics (CharMatch, Jaccard) have lower separation power for reasoning models with high paraphrase variation. The architecture is metric-agnostic: any similarity function satisfying the threshold separation property can be substituted, and semantic similarity (embedding cosine distance) is a natural extension.
Provider cooperation. Basic G2 requires access to the same model version, mitigated by negotiated replay rights, deterministic inference providers [6], or the enhanced G2 instantiation (DV-SNARK, §4.10) which reduces cooperation to one-time proof generation.
@@ 8.7 Trust Infrastructure Limitations
Non-agent node compromise. We assume non-agent nodes (CAs, humans, organizations) are uncompromised. Standard PKI mitigations apply (CT logs, TPM 2.0 trust anchors, multi-party issuance, short certificate lifetimes); our framework inherits these risks and mitigations from the PKI foundation.
Ledger operator trust. The current G3 instantiation assumes a centralized operator honest for availability and ordering; record integrity is enforced cryptographically (bilateral signatures, hash chains) regardless. Alternative instantiations (multi-operator consensus, transparency logs, blockchain anchoring) can relax the availability assumption at the cost of latency, preserving G3 semantics (§5).
Application-layer data flow. Our framework governs agent identity and capability, not the content of data flows. Complementary approaches (data-flow taint tracking, differential privacy) address this orthogonal concern.
@@ 8.8 Future Work
Formalizing the trilemma. A rigorous formalization reducing to Rice’s theorem and simulation bounds—analogous to the Gilbert–Lynch proof of CAP (2002)—would establish the trilemma as a foundational constraint on agent governance.
Hardware-attested identity. TPM 2.0 and TEE [53, 54] integration for stronger G2 guarantees without provider cooperation, addressing the trilemma by substituting compute-domain verification with trust-domain attestation.
Stealthy model substitution. Our adversarial experiment (Experiment 6) shows QLoRA-based backdoors are trivially detected. More sophisticated attacks (distillation-based backdoors, gradient-masked training) may produce stealthier substitutions; characterizing the fundamental limits of replay-based detection is an open question.
Semantic verification metrics. Embedding-based semantic similarity for reproducibility verification with formal threshold analysis, potentially extending G2’s applicability to low-determinism reasoning models.
IVC for G3. An enhanced G3 using Incrementally Verifiable Computation (e.g., Nova folding) would enable $O(1)$ audit verification regardless of chain length, deferred as current scales do not justify the ${\sim}200\times$ append overhead.
Cross-protocol governance. Extending capability binding beyond MCP/A2A to emerging agent protocols.
Dynamic trust calibration. Trust scores evolving based on ledger history, enabling governance levels to adapt automatically based on observed behavior.
@@ 9 Related Work
[…]
@@
[…]
@@
Instruction-Data Separation. A related line of work addresses the model’s inability to distinguish instructions from data within the transformer. The Instruction Hierarchy [61] trains models to prioritize system prompts over user inputs and tool outputs, establishing trust levels per input type. However, this operates at the model level via fine-tuning—it does not address the orchestration-layer problem that our capability-context separation targets: tool definitions and runtime context have different governance lifecycles (infrequent re-authorization vs. per-interaction auditing) that no amount of model training can enforce. Our separation is complementary: even a perfectly instruction-following model requires external cryptographic infrastructure to detect when its capability envelope has changed (G1), verify its computational process (G2), and produce tamper-evident interaction records (G3).
@@ 10 Conclusion
We identified silent capability escalation as a new vulnerability class in AI agent ecosystems and traced its root cause to a missing architectural distinction: the capability-context separation, which recognizes that tool definitions and runtime context have fundamentally different security semantics at the orchestration layer. From this principle, we derived three Agent Governance Requirements (G1–G3) that define what an agent ecosystem must enforce, independent of how—a governance architecture, not new cryptographic primitives.
Two structural results have prescriptive architectural consequences. The Chain Verifiability Theorem establishes that behavioral verification is a chain property: one unverifiable interior agent breaks end-to-end verification for all downstream nodes, constraining how multi-agent systems must be architected. The Bounded Divergence Theorem transforms replay-based verification into a probabilistic safety certificate, establishing software G2 as a deployable approximation of trusted execution. A multi-provider reproducibility study (9 models, 7 providers) reveals $5.8\times$ variance in inference determinism, directly connecting model characteristics to governance architecture and elevating reproducibility to a first-class model performance metric.
We validated the framework with two crypto-agnostic instantiations—basic (Ed25519, SHA-256, hash chains; 97 µs verify) and enhanced (BBS+ selective disclosure, DV-SNARK; 13.8 ms verify)—both satisfying the same nine security properties. End-to-end evaluation over 5–20 agent pipelines with real LLM calls confirms <<0.02% governance overhead and real-time detection of 7 end-to-end attack scenarios (covering G1, G2, and G3) with zero false positives, while three baselines (no governance, OAuth 2.1, OpenTelemetry) detect none. An additional 9 structural attacks are validated via automated unit tests, and 3 runtime behavioral attacks are validated via post-hoc forensic tracing.
The Agent Governance Trilemma (§8.2) reveals why no single configuration suffices: agents with unbounded capabilities (code execution, tool composition) require verification at least as expensive as the original execution, forcing an explicit tradeoff between capability, performance, and security. G1–G3 provides the infrastructure for the full governance spectrum—from compile-time safety (bounded tools, full verification) to post-hoc traceability (unbounded tools, audit-only)—making each deployment’s security posture explicit and auditable.
As the EU AI Act’s high-risk provisions take effect and NIST develops agent-specific standards, the need for cryptographic infrastructure enforcing G1–G3 will only grow. By making reproducibility a prerequisite for behavioral verifiability, our framework creates a concrete incentive for model providers to self-declare inference determinism characteristics—elevating reproducibility from an implementation detail to a first-class, security-critical model specification. We release the core mechanisms as an open-source SDK (agent-trust-kit) to enable ecosystem adoption.
Appendix A Proofs of Security Properties
This appendix contains the full proofs for all theorems stated in §6.
Proof of Property 3 (Trust Containment).
Proof of Property 4 (Capability Binding).
By collision resistance: $\Pr[H_{v}(t^{\prime})=H_{v}^{\mathit{cert}}]\leq\mathrm{negl}(\lambda)$ when $S_{v}(t^{\prime})\neq\sigma_{v}$. Forging a new certificate requires the parent’s key (same argument as Property 3). ∎∎
Proof of Property 5 (Chain Integrity).
Proof of Property 6 (Delegation Depth Enforcement).
Proof of Property 7 (Reproducibility Soundness).
Hardware G2: TEE attestation directly proves that the declared model was executed; substitution is detected with certainty. Software G2: By DiFR [1], for different models the empirical match rate is $\delta\approx 0.06$ (character-level), validated in our Experiment 5 across 9 models and 7 providers. Our empirical data shows that the fraction of cross-model output pairs exceeding any reasonable threshold $\theta$ (e.g., $\theta=0.15$) is at most $q\approx 0.02$ (i.e., $\Pr[\mathrm{CharMatch}(M(x),M^{\prime}(x))\geq\theta\mid M^{\prime}\neq M]\leq q$). Each verification prompt is therefore an independent Bernoulli trial with pass probability at most $q$. The probability that a substituted model passes all $n$ verification prompts is at most $q^{n}$, which is negligible for realistic $n$ (e.g., $q^{10}\leq 10^{-17}$). ∎∎
Proof of Theorem 3 (Bounded Divergence Under Verification).
Let $p=\Pr_{x\sim\mathcal{D}}[\mathrm{CharMatch}(M(x),M^{\prime}(x))<\theta]$ be the true failure rate. Each verification trial $Z_{i}=\mathbf{1}[\mathrm{CharMatch}(M(x_{i}),M^{\prime}(x_{i}))\geq\theta]$ is an independent Bernoulli(1p)(1-p) random variable. Observing all $n$ trials pass ($\sum Z_{i}=n$), we seek an upper confidence bound on $p$.
The probability of observing $n$ successes when the true failure rate is $p$ is $(1-p)^{n}$. The exact one-sided Clopper-Pearson upper bound at confidence level $1-\alpha$ is the value $p^{*}$ satisfying $(1-p^{*})^{n}=\alpha$, giving $p^{*}=1-\alpha^{1/n}$.
For small $p^{*}$, using the approximation $\ln(1-p)\approx-p$: from $(1-p)^{n}=\alpha$ we get $n\ln(1-p)=\ln\alpha$, hence $p\approx\ln(1/\alpha)/n$.
This is a distribution-free result: the bound depends only on the number of verification trials $n$ and the desired confidence level $\alpha$, not on the threshold $\theta$ or any distributional assumptions on the CharMatch statistic. The role of $\theta$ is solely to define what constitutes a “pass”—a higher $\theta$ increases sensitivity but may also increase false positives for the legitimate model. ∎∎
Proof of Property 8 (Credential Isolation).
Proof of Property 9 (Ledger Integrity).
Proof of Theorem 4 (Chain Verifiability).
We proceed by induction on chain position.
[…]
Proof of Property 10 (G2 Input Integrity via Bilateral Signing).
Let $v_{j}$ be a compromised agent and let $R_{i}=(\ldots,a_{\mathit{send}},a_{\mathit{recv}},\ldots,\eta_{\mathit{in}},\ldots,\mathit{sig}_{i})$ be the ledger record for the interaction delivering input to $v_{j}$, where $a_{\mathit{send}}=v_{j-1}$ and $a_{\mathit{recv}}=v_{j}$.
The record $R_{i}$ is bilaterally signed: $\mathit{sig}_{i}$ includes signatures from both $v_{j-1}$ and $v_{j}$. The input commitment $\eta_{\mathit{in}}=\mathrm{SHA\text{-}256}(\mathit{input})$ is covered by $v_{j-1}$’s signature.
Claim: $v_{j}$ cannot cause the verifier to use a fabricated input $\mathit{input}^{\prime}\neq\mathit{input}$ for replay verification.
Suppose $v_{j}$ attempts to present $\mathit{input}^{\prime}$ during forensic reconstruction. The verifier computes $\mathrm{SHA\text{-}256}(\mathit{input}^{\prime})$ and compares against $\eta_{\mathit{in}}$ in the ledger record. By collision resistance of SHA-256, $\mathrm{SHA\text{-}256}(\mathit{input}^{\prime})=\eta_{\mathit{in}}$ with negligible probability when $\mathit{input}^{\prime}\neq\mathit{input}$. Moreover, modifying $\eta_{\mathit{in}}$ in the record $R_{i}$ invalidates $v_{j-1}$’s signature (EUF-CMA of Ed25519), and is further detectable via hash-chain integrity ($h_{\mathit{prev}}$ of $R_{i+1}$ would not match). When $v_{j-1}$ is honest (Cases 1 and 2 in Property 10), the verifier independently requests $v_{j-1}$ to disclose $\mathit{input}$ and verifies it against $\eta_{\mathit{in}}$, bypassing $v_{j}$ entirely.
When $v_{j-1}$ is also compromised (Case 3), both parties may collude to present a consistent fabrication. However, Chain Verifiability (Theorem 4) already establishes $\mathrm{CVD}\leq j-1<j$: the chain’s verifiable region does not extend to $v_{j}$, so G2 replay of $v_{j}$ is not claimed to provide end-to-end guarantees in this configuration. ∎∎
Appendix B Governance Levels
The Agent Governance Trilemma (Observation 1) implies that no single governance configuration is optimal for all deployments. G1–G3 provide the infrastructure for three governance levels, each corresponding to a different position in the capability-performance-security space:
Level 3: Compile-Time Safety (high security + high performance, bounded capability). For safety-critical deployments—financial transactions, medical decisions, legal document generation—agents are restricted to bounded tools only: no code execution, no tool composition. G1 binds the complete capability envelope; G2 with full replay verification ($n\geq 200$, $\epsilon<2.3\%$ at 99% confidence) ensures behavioral fidelity; G3 provides complete audit trails. Recommended crypto backend: Enhanced G1 (BBS+ with predicate proofs, §3.5) + Enhanced G2 (DV-SNARK, §4.10).
Level 2: Sampled Verification (balanced tradeoff). For enterprise deployments—data processing, customer service, code generation—agents may use selected unbounded tools, but G2 operates in statistical mode with periodic sampling. A verification budget of $n=50$ per audit cycle bounds adversarial divergence to $\epsilon<8.9\%$ (99% confidence). G1 declares the capability envelope; G3 ensures traceability. Recommended crypto backend: Enhanced G1 (BBS+ selective disclosure) + Basic G2 (direct CharMatch).
Level 3: Post-Hoc Traceability (high capability + high performance, reduced security). For low-risk deployments—writing assistance, creative design, information retrieval—agents operate with full capability and minimal runtime overhead. G2 is set to $\mathit{none}$; G1 provides basic capability declaration; G3 provides lightweight interaction traces. Security relies on post-hoc forensic reconstruction. Recommended crypto backend: Basic G1 (Ed25519).
The governance level is declared in the agent’s certificate and recorded in the ledger, making each deployment’s security posture explicit to auditors, compliance systems, and downstream agents.